# While recognizing actions in videos, LMMs struggle to detect core interaction events

### v1

@@ 9 Detailed results
Action and object recognition.
For completeness, we also evaluated GPT-5 on the action-recognition task under the same zero-shot protocol described in Section 4 of the main manuscript. GPT-5 obtained a Top-1 accuracy of 63.6% and a perfect Top-5 accuracy of 100%, outperforming both GPT-4o and Qwen-2.5VL on this subset.
@@
We first examine a basic property: whether the model-predicted bounding box contains the true event location point. We additionally evaluated GPT-5 using the same setup, and include its results alongside GPT-4o and Qwen-2.5VL. S-Table 4 reports, for each model, the fraction of ground-truth event points that fall inside the predicted bounding box.
@@
S-Table 4: Ground-truth event point locations contained inside the model-predicted bounding boxes.
Despite this relatively high containment rate (especially for Qwen), models often predicted over-sized bounding boxes, which in many cases even exceeded the actual image dimensions. S-Table 5 presents a per-model breakdown of the Intersection-over-Union (IoU) between the predicted box and a $120\times 120$ pixel ground-truth box centered on the annotated event location.
@@
S-Table 5: Per-model IoU statistics for the event localization task.
Although the true event point often lies inside the predicted box (S-Table 4), the IoU values are extremely low (S-Table 5). This discrepancy indicates that the models tend to predict overly large bounding boxes, possibly ignoring detections of the participating objects and hands involved in the interactions. In many cases the predicted event bounding box extended beyond the frame boundaries despite explicit instructions to preserve the original spatial scale (see LABEL:supplst:bbox_prompt). Overall, all those models struggle to localize the interaction event regions reliably.
@@
S-Table 7 lists the detailed results from our experiments with (W) and without (W/O) the two modifying conditions: Reasoning and Grounding. The table includes the results under the three ICL regimes (ZS, OS, TS) for the models Qwen-2.5VL-72B and GPT-4o. The mean accuracy is evaluated for the detection of the exact true frame, where a core interaction event occurs, and also for a detection within an error tolerance of two frames off.
@@
S-Figure 5 and S-Figure 6 show additional examples of false and correct frame predictions. The examples include the models’ Chain-of-Thought, which seemingly presents a logical reasoning text for detecting an interaction event, but the grounding to the video frames is often very loose.
@@
The results of this experiment showed no improvement in the model’s performance on the test task of predicting the frame where an event occurs in the test video. The mean accuracy is reported in Tab. 6. The results suggest that the feedback session may even interfere with the main prediction task, by shifting away the model from the visual input to the number of the frame, while trying to maximize the feedback score, which is a metric on the prediction error of the frame number.
S-Table 6: Effect of feedback on the model’s performance. The true label of the second example is provided through an iterative session with feedback, indicating the gap between the predicted and the true frame. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact true frame or up to 2 frames off). The evaluation is performed under the two modifying conditions: Reasoning and Grounding. The evaluated model is Qwen-2.5VL-72B.
@@
Exact 2-off
W/O W/O $4\pm 2$ $29\pm 4$
W/O W $4\pm 2$ $29\pm 4$
W W/O $\textbf{9}\pm 2$ $\textbf{53}\pm 4$
W W $9\pm 2$ $50\pm 4$
Nevertheless, an analysis of the model’s weighted success rate and its test error shown in Fig. 7, indicate that when the model is required to provide reasoning which lead to its predictions, the test error remains low up until the 6th feedback iterations, suggesting that some learning may occur. However, as discussed in the main text, the loose grounding to the visual input, becomes even worst with this feedback approach since the focus of the model is drawn away from the image contents, trying to satisfy the feedback score around the frame number, rather than grounded visual cues.
The weighted success rate of a test trial (see Fig. 7a,b) was calculated with the conditioned probability as follows
@@ 10 Experimental Dataset
S-Table LABEL:stab:exp_dataset_details lists all 33 videos included in the experimental dataset, which was used in our experiments. For each video, we cropped short 10-frame video clips around three temporally separated core interaction events. The table includes the video ID, action template and object placeholders from the original SSv2 dataset. In addition, for each short video clip, the table includes the crop start frame, the frame where the event occurs and the type of the event, i.e., ’contact’ or ’release’. The full annotations are available online at: ssv2-contact-release-interaction-dataset.
S-Figure 5: Example false predictions of the model Qwen-2.5VL-72B. The model provides the presented chain-of-thought under the ”WITH” Reasoning condition. The examples show that the reasoning text seems logical and realistic, but the relation to the actual video frames is often very loose. A red box marks a false prediction, while the orange box marks the true frame.
S-Figure 6: Additional example predictions of the model Qwen-2.5VL-72B. The top two examples show false predictions of the model, in which the predicted frame was one frame before or after the true frame, where the event occurs. Humans can see clearly the moment of release or contact in the true frames, but the visual cues are too subtle for the models to detect. The Chain-of-Thought text lists the correct reasoning flow, but is not well grounded in the video frames. Orange and green boxes mark the true frame. Red boxes mark false predictions.
@@
S-Figure 7: Effect of feedback on test success rate and error distribution. In the TS with feedback experiment, the model goes through an iterative session, when it is provided with a feedback score until it predicts correctly the event frame of the second example. (a-b) The model’s weighted success rate (see Sec. 9) on the test video vs. the number of trials in the feedback session, without and with Reasoning, respectively. (c-d) The distribution of the test error (in frames) without and with Reasoning, respectively. The results show that when the model is not required to provide reasoning for its prediction, the feedback session does not improve the model’s performance in the test task and the test error increases with the number of feedback trials (a, c). On the other hand, when the model is required to provide reasoning, the test success rate decreases monotonically with the number of feedback trials up till the 6th iteration.
S-Table 7: Reasoning and Grounding conditions’ effect on LMMs’ performance. In the ”WITH” (W) Reasoning condition, models are instructed to describe in step-by-step the the reasoning behind their prediction of the frame where the interaction occurs. In the ”WITH” Grounding condition, prior to the main interaction detection task, models are instructed to name the target object and specify the length of the video sequence, to improve their perceptual grounding. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact or two frames off the true frame). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Qwen-2.5VL-72B and (ii) GPT-4o.
ICL Regime Reasoning Grounding Mean Accuracy (%)
Exact 2-off
Qwen-2.5VL GPT-4o Qwen-2.5VL GPT-4o
Zero Shot W/O W/O $8.1\pm 2.8$ $11.1\pm 3.2$ $57.6\pm 5.0$ $56.6\pm 5.0$
W/O W $6.1\pm 2.4$ $8.1\pm 2.8$ $46.5\pm 5.0$ $49.5\pm 5.1$
W W/O $11.1\pm 3.2$ $10.1\pm 3.0$ $54.5\pm 5.0$ $58.6\pm 5.0$
W W $10.1\pm 3.0$ $8.1\pm 2.8$ $58.6\pm 5.0$ $51.5\pm 5.0$
One Shot W/O W/O $7.2\pm 2.1$ $13.0\pm 2.2$ $56.0\pm 4.4$ $64.3\pm 3.8$
W/O W $5.2\pm 1.8$ $11.7\pm 2.2$ $51.2\pm 4.4$ $61.9\pm 4.0$
W W/O $1.6\pm 2.3 11$$1.3\pm 1.7$ $62.1\pm 3.8$ $62.4\pm 3.4$
W W $1.5\pm 2.4 11$$1.3\pm 1.7$ $60.8\pm 3.9$ $62.4\pm 3.4$
Two Shot W/O W/O $12.2\pm 2.6$ $14.7\pm 2.2$ $1.8\pm 4.6 66$$6.9\pm 3.8$
W/O W $8.6\pm 2.2$ $14.8\pm 2.4$ $58.8\pm 4.6$ $66.3\pm 3.8$
W W/O $14.0\pm 2.6$ $13.4\pm 1.8$ $4.9\pm 3.9 66$$6.6\pm 3.3$
W W $14.7\pm 2.8$ $13.5\pm 1.8$ $6.9\pm 3.8 66$$6.2\pm 3.3$
S-Table 8: Experimental dataset information.
Video ID Action template Object placeholders Event clip (10 frames)
Start frame Event frame Event type
1979 Putting [something], [something] and [something] on the table scale, eraser, sd card 12 14 release
@@
2648 Attaching [something] to [something] dummy peach, peach tree 11 18 contact
@@
3996 Putting number of [something] onto [something] books, shelf 8 13 release
@@
4042 Pushing [something] so it spins green candy 7 10 release
@@
4144 Poking [something] so that it falls over pen 12 20 contact
@@
9257 Piling [something] up kool-aid packs 8 14 release
@@
12492 Putting [something], [something] and [something] on the table keys, lock, bulb 1 7 release
@@
14990 Putting [something], [something] and [something] on the table perfume bottle, naphthalene ball, silver ring 22 25 release
@@
17127 Putting [something], [something] and [something] on the table prescribers guide book, medicine bottle, vape pen 13 17 release
@@
26039 Pushing [something] so that it falls off the table toy 1 9 contact
@@
30880 Putting [something], [something] and [something] on the table scissors, cookie cutter, grater 12 17 release
@@
41434 Stacking [number of] [something] 3, coins 0 4 contact
@@
57029 Taking [something] out of [something] tools, toolbox 0 1 contact
@@
66464 Moving [part] of [something] tuner, electric guitar 12 14 contact
@@
67618 Putting [something], [something] and [something] on the table bottle, tube, purse 0 1 release
@@
73232 Taking [something] out of [something] cd, book 4 8 contact
@@
74722 Taking [something] out of [something] phone, drawer 0 8 contact
@@
84410 Attaching [something] to [something] pen’s cover, pen 6 10 contact
@@
87327 Putting [something], [something] and [something] on the table grater, whisk, corkscrew 4 11 release
@@
92626 Poking [something] so that it spins around flashlight 2 7 contact
@@
95238 Attaching [something] to [something] toy train engine, its coach 3 7 contact
@@
96903 Rolling [something] on a flat surface perfume 2 7 contact
@@
153413 Putting [something], [something] and [something] on the table fork, spoon, dish 10 17 release
@@
158080 Putting [something], [something] and [something] on the table toothpick container, showpiece, padlock 2 6 release
@@
158915 Putting [something], [something] and [something] on the table mug, spoon, gum 1 5 release
@@
163090 Putting [something], [something] and [something] on the table popcorn, vicks vaporub bottle, purple water bottle 5 10 release
@@
164784 Pushing [something] so that it almost falls off but doesn’t roll 3 6 contact
@@
166894 Poking a stack of [something] without the stack collapsing lincoln logs 15 19 contact
@@
175159 Stacking [number of] [something] 5, hot pads 6 9 contact
@@
175167 Piling [something] up water color containers 4 8 contact
@@
181367 Piling [something] up shoes 0 2 contact
@@
186500 Pulling [something] onto [something] nail clipper, envelope 11 17 contact
@@
217743 Putting [something], [something] and [something] on the table glass vase, child’s shoe, coffee mug 21 24 release

### v2

@@ 9 Detailed results
We first examine a basic property: whether the model-predicted bounding box contains the true event location point. Tab. S1 reports, for each model, the fraction of ground-truth event points that fall inside the predicted bounding box.
@@
Table S1: Ground-truth event point locations contained inside the model-predicted bounding boxes.
Although the true event point often lies inside the predicted box (Tab. S1), the Intersection-over-Union (IoU) between the predicted box and a $120\times 120$ pixels box around the true location point are extremely low (Tab. S2, Fig. S4). This discrepancy indicates that while the models often detect the hand or the target object, they tend to ignore important image regions around the target objects and hands, which contain critical information about the interactions. In some cases the predicted bounding box extended beyond the frame boundaries despite explicit instructions to preserve the original spatial scale (see LABEL:supplst:bbox_prompt). Overall, the models struggle to localize the interaction event regions reliably.
@@
Table S2: Per-model IoU statistics for the event localization task.
@@
Tab. S3 and Tab. S4 complements the results for the models Qwen-2.5VL-72B and GPT-4o from our experiments on the two modifying conditions: Reasoning and Grounding.
Table S3: Reasoning effect on LMMs’ performance. Models are instructed to describe step-by-step the reasoning behind their prediction of the frame where the interaction occurs. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact or one-frame off the true frame). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Qwen-2.5VL-72B and (ii) GPT-4o.
ICL RSN Mean Accuracy Percentage (SD)
Exact 1-off
Qwen GPT-4o Qwen GPT-4o
ZS W/O $7.07$ $(1.83)$ $9.60$ $(2.10)$ $31.31$ $(3.30)$ $28.79$ $(3.23)$
W 10.61 $(2.19)$ $9.09$ $(2.10)$ $28.79$ $(3.23)$ 31.82 $(3.32)$
OS W/O $6.21$ $(1.38)$ 12.34 $(1.55)$ $29.81$ $(2.81)$ 36.97 $(2.63)$
W $11.53$ $(1.66)$ $11.29$ $(1.20)$ $36.12$ $(2.58)$ $36.62$ $(2.36)$
TS W/O $10.39$ $(1.72)$ 14.72 $(1.63)$ $36.44$ $(3.05)$ $38.61$ $(2.86)$
W $14.34$ $(1.90)$ $13.44$ $(1.29)$ $39.31$ $(2.76)$ 40.52 $(2.44)$
Table S4: Grounding effect on LMMs’ performance. Prior to the main interaction detection task, models are instructed to name the target object and specify the length of the video sequence, to improve their perceptual grounding. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact true frame or up to one- frame off). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Qwen-2.5VL-72B and (ii) GPT-4o.
ICL GRND Mean Accuracy Percentage (SD)
Exact 1-off
Qwen GPT-4o Qwen GPT-4o
ZS W/O $9.60$ $(2.10)$ 10.61 $(2.19)$ $30.30$ $(3.27)$ 32.32 $(3.33)$
W $8.08$ $(1.94)$ $8.08$ $(1.94)$ $29.80$ $(3.26)$ $28.28$ $(3.21)$
OS W/O $9.40$ $(1.58)$ 12.13 $(1.39)$ $34.59$ $(2.77)$ 37.29 $(1.39)$
W $8.34$ $(1.50)$ $11.49$ $(1.39)$ $31.34$ $(2.63)$ $36.49$ $(3.31)$
TS W/O $13.11$ $(1.72)$ $14.04$ $(1.44)$ $38.92$ $(2.94)$ 42.06 $(2.57)$
W $11.63$ $(1.77)$ 14.11 $(1.49)$ $36.83$ $(2.89)$ $37.07$ $(2.73)$
@@
Fig. S2 and Fig. S3 show additional examples of false and correct frame predictions. The examples include the models’ Chain-of-Thought, which seemingly presents a logical reasoning text for detecting an interaction event, but the grounding to the video frames is often very loose.
@@
The results of this experiment showed no improvement in the model’s performance on the test task of predicting the frame where an event occurs in the test video. The mean accuracy is reported in Tab. S5. The results suggest that the feedback session may even interfere with the main prediction task, by shifting away the model from the visual input to the number of the frame, while trying to maximize the feedback score, which is a metric on the prediction error of the frame number.
Table S5: Effect of feedback on the model’s performance. The true label of the second example is provided through an iterative session with feedback, indicating the gap between the predicted and the true frame. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact true frame or up to one-frame off). The evaluation is performed under the two modifying conditions: Reasoning and Grounding. The evaluated model is Qwen-2.5VL-72B.
@@
Exact 1-off
W/O W/O $2.55\pm 1.12$ $18.37\pm 3.43$
W/O W $5.61\pm 1.76$ $21.94\pm 3.70$
W W/O $12.59\pm 2.63$ $\textbf{41.84}\pm 4.15$
W W $\textbf{16.33}\pm 3.06$ $39.80\pm 4.10$
Nevertheless, an analysis of the model’s weighted success rate and its test error shown in Fig. S6, indicate that when the model is required to provide reasoning which lead to its predictions, the test error remains low up until the 6th feedback iterations, suggesting that some learning may occur. However, as discussed in the main text, the loose grounding to the visual input, becomes even worst with this feedback approach since the focus of the model is drawn away from the image contents, trying to satisfy the feedback score around the frame number, rather than grounded visual cues.
The weighted success rate of a test trial (see Fig. S6a,b) was calculated with the conditioned probability as follows
@@
It should be noted that we also tested a setting in where the images were fed to the model in an arbitrary order. This change in image ordering alone resulted in a significant drop in performance, as seen in Tab. S5, despite an explicit instruction that was given to the model that the frames should be treated in consecutive order. From this experiment we conclude that in the functioning of LMMs there is no notion of ordering of frames, unless they are fed in as an ordered sequence.
Chain-of-Thought (CoT) Tuning.
In this ICL few-shot experiment, examples included specific CoT descriptions for detecting the frame where the contact/release events occurred. The results in Tab. S6 show that the accuracy does not improve beyond 2 examples. However, introducing an explicit CoT description in the examples yielded enhanced accuracy compared to the experiment without CoT (see Tab. S3).
Table S6: Results for the CoT experiment expressed in mean accuracy within an error tolerance of one-frame off from the true frame in each test sample. The evaluated model is Qwen-2.5VL-72B.
Number of Examples Mean Accuracy (%)
1 $51.52\pm 5.02$
2 $55.56\pm 4.99$
3 $45.45\pm 5.00$
4 $44.44\pm 4.99$
5 $39.39\pm 4.91$
6 $43.43\pm 4.98$
7 $38.38\pm 4.89$
8 $48.48\pm 5.02$
(a) Qwen-2.5VL-72B
(b) GPT-4o
Figure S1: Mean accuracy vs. detection error tolerance. A correct detection of the models represents a predicted frame within the allowed error tolerance, where an error tolerance of zero means the exact true frame was predicted. Results of Qwen-2.5VL-72B (a) and GPT-4o (b) are shown for the difference ICL regimes under the ”with reasoning” condition. Note, that the length of all videos in the experimental dataset is 10 frames.
@@ 10 Experimental Dataset
Our experimental dataset included 33 videos from SSv2. For each video, we cropped short 10-frame video clips around three temporally separated core interaction events, resulting with 99 event clips. By construction, event frames are evenly distributed over the 10-frame window to avoid any bias. In addition, we have ensured that only a single event appears in the time window. Tab. LABEL:stab:exp_dataset_details includes the video ID, action template and object placeholders from the original SSv2 dataset. In addition, for each short video clip, the table includes the crop start frame, the frame where the event occurs and the type of the event, i.e., ’contact’ or ’release’. All video clips in this set are at 12 fps. Fig. S5 presents a few examples of the 10-frames clips and annotated event frames used in the evaluation. The full annotations are available online at: ssv2-contact-release-interaction-dataset.
Figure S2: Example false predictions. The model (here Qwen-2.5VL-72B) provides the presented chain-of-thought under the ”WITH” Reasoning condition. The examples show that the reasoning text seems logical and realistic, but the relation to the actual video frames is often very loose. A red box marks a false prediction, while the orange box marks the true frame.
Figure S3: Additional example predictions. The top two examples show false predictions of the model (here Qwen-2.5VL-72B), in which the predicted frame was one frame before or after the true frame, where the event occurs. Humans can see clearly the moment of release or contact in the true frames, but the visual cues are too subtle for the models to detect. The Chain-of-Thought text lists the correct reasoning flow, but is not well grounded in the video frames. Orange and green boxes mark the true frame. Red boxes mark false predictions.
Figure S4: Bounding box event detection examples of Gemini-3-Pro model. Upper two rows: Failure detection examples. Bottom row: Successful detection examples. Green dot and bbox indicate the ground truth event location and a $120\times 120$ pixels bboux around it. Red bbox indicates the model’s prediction.
Figure S5: Full interaction sequence examples from the evaluation dataset. A green box marks the labeled event frame.
@@
Figure S6: Effect of feedback on test success rate and error distribution. In the TS with feedback experiment, the model goes through an iterative session, when it is provided with a feedback score until it predicts correctly the event frame of the second example. (a-b) The model’s weighted success rate (see Sec. 9) on the test video vs. the number of trials in the feedback session, without and with Reasoning, respectively. (c-d) The distribution of the test error (in frames) without and with Reasoning, respectively. The results show that when the model is not required to provide reasoning for its prediction, the feedback session does not improve the model’s performance in the test task and the test error increases with the number of feedback trials (a, c). On the other hand, when the model is required to provide reasoning, the test success rate decreases monotonically with the number of feedback trials up till the 6th iteration.
Table S7: Experimental dataset information.
Video ID Action template Object placeholders Frame size Event clip [^1] All video clips in this set are at 12 fps. (10 frames)
(width,height) Start frame Event frame Event type
1979 Putting [something], [something] and [something] on the table scale, eraser, sd card $(320,240)$ 12 14 release
@@
2648 Attaching [something] to [something] dummy peach, peach tree $(427,240)$ 11 18 contact
@@
3996 Putting number of [something] onto [something] books, shelf $(427,240)$ 8 13 release
@@
4042 Pushing [something] so it spins green candy $(427,240)$ 7 10 release
@@
4144 Poking [something] so that it falls over pen $(320,240)$ 12 20 contact
@@
9257 Piling [something] up kool-aid packs $(320,240)$ 8 14 release
@@
12492 Putting [something], [something] and [something] on the table keys, lock, bulb $(293,240)$ 1 7 release
@@
14990 Putting [something], [something] and [something] on the table perfume bottle, naphthalene ball, silver ring $(320,240)$ 22 25 release
@@
17127 Putting [something], [something] and [something] on the table prescribers guide book, medicine bottle, vape pen $(432,240)$ 13 17 release
@@
26039 Pushing [something] so that it falls off the table toy $(427,240)$ 1 9 contact
@@
30880 Putting [something], [something] and [something] on the table scissors, cookie cutter, grater $(427,240)$ 12 17 release
@@
41434 Stacking [number of] [something] 3, coins $(293,240)$ 0 4 contact
@@
57029 Taking [something] out of [something] tools, toolbox $(427,240)$ 0 1 contact
@@
66464 Moving [part] of [something] tuner, electric guitar $(320,240)$ 12 14 contact
@@
67618 Putting [something], [something] and [something] on the table bottle, tube, purse $(427,240)$ 0 1 release
@@
73232 Taking [something] out of [something] cd, book $(427,240)$ 4 8 contact
@@
74722 Taking [something] out of [something] phone, drawer $(427,240)$ 0 8 contact
@@
84410 Attaching [something] to [something] pen’s cover, pen $(427,240)$ 6 10 contact
@@
87327 Putting [something], [something] and [something] on the table grater, whisk, corkscrew $(427,240)$ 4 11 release
@@
92626 Poking [something] so that it spins around flashlight $(427,240)$ 2 7 contact
@@
95238 Attaching [something] to [something] toy train engine, its coach $(427,240)$ 3 7 contact
@@
96903 Rolling [something] on a flat surface perfume $(320,240)$ 2 7 contact
@@
153413 Putting [something], [something] and [something] on the table fork, spoon, dish $(427,240)$ 10 17 release
@@
158080 Putting [something], [something] and [something] on the table toothpick container, showpiece, padlock $(427,240)$ 2 6 release
@@
158915 Putting [something], [something] and [something] on the table mug, spoon, gum $(427,240)$ 1 5 release
@@
163090 Putting [something], [something] and [something] on the table popcorn, vicks vaporub bottle, purple water bottle $(427,240)$ 5 10 release
@@
164784 Pushing [something] so that it almost falls off but doesn’t roll $(320,240)$ 3 6 contact
@@
166894 Poking a stack of [something] without the stack collapsing lincoln logs $(427,240)$ 15 19 contact
@@
175159 Stacking [number of] [something] 5, hot pads $(427,240)$ 6 9 contact
@@
175167 Piling [something] up water color containers $(427,240)$ 4 8 contact
@@
181367 Piling [something] up shoes $(320,240)$ 0 2 contact
@@
186500 Pulling [something] onto [something] nail clipper, envelope $(427,240)$ 11 17 contact
@@
217743 Putting [something], [something] and [something] on the table glass vase, child’s shoe, coffee mug $(427,240)$ 21 24 release